# Factorized Tensor Networks for Multi-Task and Multi-Domain Learning

### v1

@@ 4.2 Multi-task dense prediction
Dataset. The widely-used NYUD dataset [73] with 795 training and 654 testing images of indoor scenes is used for dense prediction experiments in multi-task learning. The dataset contains four tasks: edge detection (Edge), semantic segmentation (SemSeg), surface normals estimation (Normals), and depth estimation (Depth). We follow the same data-augmentation technique as used by [9].
@@
Training details. ResNet-18 is used as the backbone network, and DeepLabv3+ [75] as the decoder architecture. The Fine-Tuning and Feature-Extractor experiments are implemented in the same way as in the classification-based experiments above. We showed experiments for FTNs with $R\in\{1,10,20,30\}$. Further details are in the supplementary material.
Results. Table 4 shows the performance of FTN with various ranks and of other baseline comparison methods for dense prediction tasks on the NYUD dataset. We observe performance improvement by increasing flexibility through higher rank. FTN with rank-30 performs better than all comparison methods and utilizes the least number of parameters. Also, on the Depth and Edge task we can attain good performance by using only rank-20. We take the performance of baseline comparison methods from the RCM paper [9] as we run our experiments under the same setting.
Table 4: Dense prediction performance on NYUD dataset using ResNet-18 backbone with DeepLabv3+ decoder. The proposed FTN approach with $R=\{1,10,20,30\}$ and other methods. The best performing method in bold.
Methods Params (Abs) Semsegnormal-$\uparrow$ Depthnormal-$\downarrow$ Normalsnormal-$\downarrow$ Edgenormal-$\uparrow$
Single Task 4×4$\times$4 × (44.68M) 35.34 0.56 22.20 73.5
Decoder only 1×1$\times$1 × (11.17M) 24.84 0.71 28.56 71.3
Decoder + BN only 1.002×1.002$\times$1.002 × (11.19M) 29.26 0.61 24.82 71.3
ASTMT (R-18 w/o SE) [10] 1.25×1.25$\times$1.25 × (13.99M) 30.69 0.60 23.94 68.60
ASTMT (R-26 w SE) [10] 2.00×2.$0$ (22.42M) 30.07 0.63 24.32 73.50
Series RA [3] 1.56×1.56$\times$1.56 × (17.51M) 31.87 0.60 23.35 67.56
Parallel RA [6] 1.50×1.50$\times$1.50 × (16.77M) 32.13 0.59 23.20 68.02
RCM [9] 1.56×1.56$\times$1.56 × (17.49M) 34.20 0.57 22.41 68.44
@@
FTN, R=10 1.03×1.03$\times$1.03 × (11.54M) 33.66 0.57 22.15 73.5
FTN, R=20 1.06×1.06$\times$1.06 × (11.89M) 34.06 0.55 21.84 73.9
FTN, R=30 1.09×1.09$\times$1.09 × (12.24M) 35.46 0.56 21.78 73.8
@@ 5 Conclusion
We have proposed a simple, parameter-efficient, architecture-agnostic, and easy-to-implement FTN method that adapts to new unseen domains/tasks using low-rank task-specific tensors. In our work, we have shown FTN requires the least number of parameters compared to other baseline methods in MDL/MTL experiments and attains better or comparable performance. We can adapt the backbone network in a flexible manner by adjusting the rank according to the complexity of the domain/task. We conducted experiments with different convolutional backbones and transformer architectures for various datasets to demonstrate that FTN outperforms existing methods.
Acknowledgements
This work is supported in part by AFOSR award FA9550-21-1-0330 and ONR award N00014-19-1-2264.
@@ Acknowledgements
In this section, we present additional material and details to supplement the main paper.
Appendix A Dataset and training details
A.1 Imagenet-to-sketch dataset
@@
We additionally compare the proposed FTN under differnet backbone architecture, EfficientNet-B4. I Table 5 we have shown the performance for Fine-Tuning, Feature-Extractor, and FTN with for $R\in\{1,10\}$. FTN irrespective of backbone architecture is giving performance close to Fine-Tuning.
Table 5: Performance on ImageNet-to-Sketch dataset under incremental setting using EfficientNet-B4 backbone with our FTN approach.
@@ A.2 DomainNet dataset
Table 6: Performance on DomainNet dataset under joint setting using resnet-34 backbone (initialized with jointly trained weights) with our FTN approach along with comparison methods.
Methods Params (Abs) Clipart Sketch Paint Quick Info Real mean
Fine-tuning 6×6$\times$6 × (127.68M) 77.43 69.25 69.21 71.61 41.50 80.74 68.29
Feature extractor 1×1$\times$1 × (21.28M) 76.67 65.2 65.26 52.97 35.05 76.08 61.87
FC and BN only 1.004×1.004$\times$1.004 × (21.35M) 77.07 68.34 68.76 69.06 40.63 79.07 67.15
Adashare 1×1$\times$1 × (21.28M) 75.88 63.96 67.90 61.17 31.52 76.90 62.88
TAPS 1.46×1.46$\times$1.46 × (31.06M) 76.98 67.81 67.91 70.18 39.30 78.91 66.84
FTN, R=1 1.008×1.008$\times$1.008 × (21.45M) 77.13 68.10 68.50 69.41 40.04 79.49 67.11
A.3 NYUD dataset
@@
A.4 Visual Decathlon Challenge dataset
We showed the performance of FTN on the visual decathlon challenge dataset. The experiments are performed on resnet-18 backbone with pre-trained weights initialized from Imagenet. In Table 7, we report the performance using FTN with rank R=1, 20 and comparison with Fine-Tuning and Feature-Extractor. We also report the total number of parameters used by each method. The FTN with rank, R=20 achieves accuracy significantly better than Feature-Extractor (and comparable to Fine-Tuning networks) while using only $\sim~{}6\%$ additional parameters compared to Feature-Extractor.
Table 7: Performance on Visual Decathlon Challenge dataset using resnet-18 backbone with FTN approach using R=1,20 along with the number of parameters.
@@ A.4 Visual Decathlon Challenge dataset
Appendix B Effect on performance with different number of low-rank factors.
@@
We performed an experiment by removing the low-rank factors from our trained FTN backbone network at different thresholds. We perform this experiment on five domains of the Imagenet-to-sketch dataset and compute the $\ell_{2}$-norm of $\Delta\mathbf{W}$ at every layer. We selected equally spaced threshold values from the minimum and maximum $\ell_{2}$-norm and removed the low-rank factors below the threshold. The performance vs. the number of parameters of low-rank layers for different thresholds is shown in Figure 3. We observe a drop in performance on every domain as we increase the threshold and reduce the number of layers from the backbone. Interestingly, when we reduce the number of layers from 52 to 28 on the Flowers and CUBS dataset, we did not see a significant drop in accuracy.
@@
Figure 3: Performance on five domains of the Imagenet-to-sketch dataset as we remove the low-rank parameters. We selected the number of layers in the backbone based on a moving threshold. We annotate the specified threshold at each marker point and the number of affected layers (in parentheses).
Appendix C Visualization of changes in FTN with low-rank factors
@@
We present the norm of low-rank factors at every adapted layer in the backbone of our FTN as a heatmap in Figures 45. The colors indicate relative norms because we normalized them for every network in the range 0 to 1 to highlight the relative differences. Figure 4 presents results on five domains of the Imagenet-to-sketch dataset (resnet-50 backbone), adapting every layer with rank-50 FTN. We observe the maximum changes in the last layer of the backbone network instead of the initial layers. We also show a similar trend on the DomainNet dataset with resnet-34 backbone where maximum changes occur in the network’s later layer (see Figure 6). We observe from Figure 5 that on the wikiart domain of the Imagenet-to-sketch dataset, the layers in the backbone network become more adaptive upon increasing the rank of FTN. The rank-50 FTN has more task adaptive layers than the rank-1 FTN on the wikiart domain.
@@
Figure 4: Norm of low-rank factors in the adapted backbone layers for different domains of the Imagenet-to-sketch dataset with $R=50$.
@@
Figure 5: Norm of low-rank factors in the adapted backbone layers for different values of $R\in\{1,5,10,15,20,25,50\}$ with the wikiart domain of the Imagenet-to-sketch dataset.
@@
Figure 6: Norm of low-rank factors in the adapted backbone layers for different domains of the DomainNet dataset with $R=40$.
Appendix D Results on DomainNet dataset under joint setting
@@
We performed additional experiments under a joint learning setup on the DomainNet dataset. A single model is trained jointly on all the domains of the dataset with Fine-Tuning setup. Table 6 summarizes the results for the performance of the DomainNet dataset under a joint setup. The domains under this setting share information among each other for all parameters. The Fine-Tuning experiment achieves the overall best performance, but at the expense of a large number of parameters. We observe poor performance for the shared Feature Extractor, since that does not learn any additional task-specific parameters. The results in the third row show that by changing just the task-specific batchnorm layers in the jointly trained backbone, we can achieve better results than TAPS and Adashare. Our FTN with just $R=1$ also outperforms Adashare and TAPS.
Appendix E Analysis on $\Delta W_{t}$ in FTN vs post-training factorization
@@
We have done experimental results to show that FTN learns explicit low-rank representation instead of a module that stores low-rank tensors. For post-training factorization, we consider $\Delta W_{t}=W_{t}-W$, i.e. computing the difference between the weights of the fine-tuned network ($W_{t}$) and the original network ($W$). We have compressed or performed low-rank approximation on the trained $\Delta W_{t}$ using CP, Tensor train (TT), and SVD matrix factorization. We performed experiments on ImageNet-to-Sketch dataset with resnet-50 and efficient-net backbone with $R\in\{1,10,20,30,40,50\}$. Table 8 depicts that low-rank approximation of trained $\Delta W_{t}$ causes severe performance degradation. For instance, R=1, average accuracy for ImageNet-to-Sketch with resnet-50 drops to nearly 27%. As R increases, the accuracy improves for all the factorization approaches, but they remain lower than FTN results. From Table 10 the difference between FTN and post-training factorization is even more striking in the case of Efficient-Net. For instance, R=1, average accuracy drops to nearly 5%. As R increases, the number of parameters increases sharply but even with 2x more parameters, the average accuracy of post-training low-rank factorization remains lower than FTN with R=1. The Table 9, 11 depicts the approximation error (averaged over all the layers) for different values R and every task. We report average error and standard deviation over the layers for a subset of R. We observe that as R increases approximation error reduces, which leads to improvement in accuracy. Nevertheless, approximation error remains significant even with large R, which is the reason why accuracy remains significantly less than fine-tuned weights. In all cases, accuracy for tasks depends on approximation error (Flowers and CUBS do better than Wikiart and Sketch because the approximation error is small).
@@
Table 8: Low-rank factorization of weight increments on ImageNet-to-Sketch dataset with resnet-50 backbone.
@@
Fine-Tuning(arbitrary $\Delta W_{t}$ = $W_{t}$$W$) 6x(141M) 95.69 78.42 81.02 91.44 83.32 85.98
@@
Table 9: Average (std deviation) approximation error for low-rank factorization of weight increments $\Delta W$ on ImageNet-to-Sketch dataset with resnet-50 backbone.
@@
Table 10: Low-rank factorization of weight increments on ImageNet-to-Sketch dataset with efficient-net backbone.
@@
Fine-Tuning (arbitrary $\Delta W_{t}$ = $W_{t}$$W$) 6x(105.24M) 96.08 78.72 80.9 92.81 83.67 86.43
@@
Table 11: Average (std deviation) approximation error for low-rank factorization of weight increments $\Delta W$ on ImageNet-to-Sketch dataset with efficient-net backbone.
@@
Appendix F Multi-domain image generation
@@
Dataset. We used the multi-domain Transient Attributes [79] dataset. The dataset contains a total of 8571 images with 40 annotated attribute labels. Each label is associated with a score in the $[-1,1]$ range. We utilize the associated confidence score for each season to build our collection of images for each season. Additionally, we only selected images that were captured during daytime. Our training set consisted of $1875$ summer, $1405$ spring, $1353$ autumn, and $2566$ winter images. We normalized each image to the range of $[-1,1]$ and resized them to a resolution of $128\times 128$.
@@
Figure 7: Generated images for different seasons using FTN.
Table 12: Image generation PSNR for different methods and seasons
@@
Table 13: FTN PSNR for image generation under $R=\{20,50\}$
@@
Results. Table 12 shows the performance, in terms of Peak Signal-to-Noise Ratio (PSNR), for three methods. Our proposed FTN network achieved a comparable performance to the single-domain networks. Each single domain network has $71.4$M trainable parameters, while the FTN network adds an additional $3.9$M parameters per domain over the base network. Figure 7 and 8 show examples of images generated by our proposed FTN network. In addition, table 13 shows the average performance of our FTN network under different rank settings. We observe a performance increase by increasing the rank $R$ of our low-rank factors.
Figure 8: Additional generated images for different seasons using FTN.

### v2

@@ 4.2 Multi-task dense prediction
Dataset. The widely-used NYUD dataset with 795 training and 654 testing images of indoor scenes is used for dense prediction experiments in multi-task learning. The dataset contains four tasks: edge detection (Edge), semantic segmentation (SemSeg), surface normals estimation (Normals), and depth estimation (Depth). We follow the same data-augmentation technique as used by [16].
@@
Training details. ResNet-18 is used as the backbone network, and DeepLabv3+ as the decoder architecture. The Fine-Tuning and Feature-Extractor experiments are implemented in the same way as in the classification-based experiments above. We showed experiments for FTNs with $R\in\{1,10,20,30\}$. Further details are in the supplementary material.
Results. Table 4 shows the performance of FTN with various ranks and of other baseline comparison methods for dense prediction tasks on the NYUD dataset. We observe performance improvement by increasing flexibility through higher rank. FTN with rank-30 performs better than all comparison methods and utilizes the least number of parameters. Also, we attain good performance on the Depth and Edge task by using only rank-20. We take the performance of baseline comparison methods from the RCM paper [16] as we run our experiments under the same setting.
Section 6 of the supplementary materials presents additional experiments on the multi-domain image generation application using the FTN method.
Table 4: Dense prediction performance on NYUD dataset using ResNet-18 backbone with DeepLabv3+ decoder. The proposed FTN approach with $R=\{1,10,20,30\}$ and other methods. Bold and underline indicate the best and second-best results, respectively.
@@
Single Task $4\times$ 35.34 0.56 22.20 73.5
Decoder only $1\times$ 24.84 0.71 28.56 71.3
Decoder + BN only $1.002\times$ 29.26 0.61 24.82 71.3
ASTMT (R-18) [27] $1.25\times$ 30.69 0.60 23.94 68.60
ASTMT (R-26+SE) [27] $2.00\times$ 30.07 0.63 24.32 73.50
Series RA [37] $1.56\times$ 31.87 0.60 23.35 67.56
Parallel RA [37] $1.50\times$ 32.13 0.59 23.20 68.02
RCM [16] $1.56\times$ 34.20 0.57 22.41 68.44
FTN, R=1 $1.005\times$ 29.83 0.60 23.56 72.7
FTN, R=10 $1.03\times$ 33.66 0.57 22.15 73.5
FTN, R=20 $1.06\times$ 34.06 0.55 21.84 73.9
FTN, R=30 $1.09\times$ 35.46 0.56 21.78 73.8
@@ 5 Conclusion
We have proposed a simple, parameter-efficient, architecture-agnostic, and easy-to-implement FTN method that adapts to new unseen domains/tasks using low-rank task-specific tensors. Our work shows that FTN requires the least number of parameters compared to other baseline methods in MDL/MTL experiments and attains better or comparable performance. We can adapt the backbone network in a flexible manner by adjusting the rank according to the complexity of the domain/task. We conducted experiments with different convolutional backbones and transformer architectures for various datasets to demonstrate that FTN outperforms existing methods.
Future work. In our current approach, we used a fixed rank for each layer. Moving forward, we can explore adaptively selecting the rank for different layers, which may further reduce the number of parameters. MDL/MTL models are often challenged by task interference or negative transfer learning when conflicting tasks are trained together. Future work can address this by investigating which tasks or domains should be learned jointly to mitigate such drawbacks. Additionally, while our method requires a separate forward pass for each task due to the shared backbone, we could further explore branched or tree-structured models that enable task-specific layer sharing to reduce latency.
@@
S1 Dataset and training details
S1.1 Imagenet-to-sketch dataset
@@
We additionally compare the proposed FTN under differnet backbone architecture, EfficientNet-B4. I Table S5 we have shown the performance for Fine-Tuning, Feature-Extractor, and FTN with for $R\in\{1,10\}$. FTN irrespective of backbone architecture is giving performance close to Fine-Tuning.
Table S5: Performance on ImageNet-to-Sketch dataset under incremental setting using EfficientNet-B4 backbone with our FTN approach.
@@
S1.3 NYUD dataset
@@
S1.4 Visual Decathlon Challenge dataset
We showed the performance of FTN on the visual decathlon challenge dataset. The experiments are performed on resnet-18 backbone with pre-trained weights initialized from Imagenet. In Table S6, we report the performance using FTN with rank R=1, 20 and comparison with Fine-Tuning and Feature-Extractor. We also report the total number of parameters used by each method. The FTN with rank, R=20 achieves accuracy significantly better than Feature-Extractor (and comparable to Fine-Tuning networks) while using only $\sim~6\%$ additional parameters compared to Feature-Extractor.
Table S6: Performance on Visual Decathlon Challenge dataset using resnet-18 backbone with FTN approach using R=1,20 along with the number of parameters.
@@
S2 Effect on performance with different number of low-rank factors.
@@
We performed an experiment by removing the low-rank factors from our trained FTN backbone network at different thresholds. We perform this experiment on five domains of the Imagenet-to-sketch dataset and compute the $\ell_{2}$-norm of $\Delta\mathbf{W}$ at every layer. We selected equally spaced threshold values from the minimum and maximum $\ell_{2}$-norm and removed the low-rank factors below the threshold. The performance vs. the number of parameters of low-rank layers for different thresholds is shown in Figure S3. We observe a drop in performance on every domain as we increase the threshold and reduce the number of layers from the backbone. Interestingly, when we reduce the number of layers from 52 to 28 on the Flowers and CUBS dataset, we did not see a significant drop in accuracy.
@@
Figure S3: Performance on five domains of the Imagenet-to-sketch dataset as we remove the low-rank parameters. We selected the number of layers in the backbone based on a moving threshold. We annotate the specified threshold at each marker point and the number of affected layers (in parentheses).
S3 Visualization of changes in FTN with low-rank factors
@@
We present the norm of low-rank factors at every adapted layer in the backbone of our FTN as a heatmap in Figures S4S5. The colors indicate relative norms because we normalized them for every network in the range 0 to 1 to highlight the relative differences. Figure S4 presents results on five domains of the Imagenet-to-sketch dataset (resnet-50 backbone), adapting every layer with rank-50 FTN. We observe the maximum changes in the last layer of the backbone network instead of the initial layers. We also show a similar trend on the DomainNet dataset with resnet-34 backbone where maximum changes occur in the network’s later layer (see Figure S6). We observe from Figure S5 that on the wikiart domain of the Imagenet-to-sketch dataset, the layers in the backbone network become more adaptive upon increasing the rank of FTN. The rank-50 FTN has more task adaptive layers than the rank-1 FTN on the wikiart domain.
@@
Figure S4: Norm of low-rank factors in the adapted backbone layers for different domains of the Imagenet-to-sketch dataset with $R=50$.
@@
Figure S5: Norm of low-rank factors in the adapted backbone layers for different values of $R\in\{1,5,10,15,20,25,50\}$ with the wikiart domain of the Imagenet-to-sketch dataset.
@@
Figure S6: Norm of low-rank factors in the adapted backbone layers for different domains of the DomainNet dataset with $R=40$.
Table S7: Performance on DomainNet dataset under joint setting using resnet-34 backbone (initialized with jointly trained weights) with our FTN approach along with comparison methods.
Methods Params (Abs) Clipart Sketch Paint Quick Info Real mean
Fine-tuning $6\times$ (127.68M) 77.43 69.25 69.21 71.61 41.50 80.74 68.29
Feature extractor(unified head) $1\times$ (21.28M) 33.74 73.97 73.13 54.75 39.49 63.64 56.45
Feature extractor $1\times$ (21.28M) 76.67 65.2 65.26 52.97 35.05 76.08 61.87
FC and BN only $1.004\times$ (21.35M) 77.07 68.34 68.76 69.06 40.63 79.07 67.15
Adashare [40] $1\times$ (21.28M) 75.88 63.96 67.90 61.17 31.52 76.90 62.88
TAPS [45] $1.46\times$ (31.06M) 76.98 67.81 67.91 70.18 39.30 78.91 66.84
FTN, R=1 $1.008\times$ (21.45M) 77.13 68.10 68.50 69.41 40.04 79.49 67.11
S4 Results on DomainNet dataset under joint setting
@@
We performed additional experiments under a joint learning setup on the DomainNet dataset. A single model is trained jointly on all the domains of the dataset with Fine-Tuning setup. Table S7 summarizes the results for the performance of the DomainNet dataset under a joint setup. The domains under this setting share information among each other for all parameters. The Fine-Tuning experiment achieves the overall best performance, but at the expense of a large number of parameters. We observe poor performance for the shared Feature Extractor, since that does not learn any additional task-specific parameters. For Feature Extractor with unified head, we combined all six domains into a single dataset and then trained a network that uses a single backbone and classifier. From the second row, we observe significant performance improvements in domains such as sketch and painting using the unified classifier. However, in other domains like clipart, quickgraph, and real, the accuracy was lower than when models were trained independently for each domain. The results in the fourth row show that by changing just the task-specific batchnorm layers in the jointly trained backbone, we can achieve better results than TAPS and Adashare. Our FTN with just $R=1$ also outperforms Adashare and TAPS.
S5 Analysis on $\Delta W_{t}$ in FTN vs post-training factorization
@@
We have done experimental results to show that FTN learns explicit low-rank representation instead of a module that stores low-rank tensors. For post-training factorization, we consider $\Delta W_{t}=W_{t}-W$, i.e. computing the difference between the weights of the fine-tuned network ($W_{t}$) and the original network ($W$). We have compressed or performed low-rank approximation on the trained $\Delta W_{t}$ using CP, Tensor train (TT), and SVD matrix factorization. We performed experiments on ImageNet-to-Sketch dataset with resnet-50 and efficient-net backbone with $R\in\{1,10,20,30,40,50\}$. Table S8 depicts that low-rank approximation of trained $\Delta W_{t}$ causes severe performance degradation. For instance, R=1, average accuracy for ImageNet-to-Sketch with resnet-50 drops to nearly 27%. As R increases, the accuracy improves for all the factorization approaches, but they remain lower than FTN results. From Table S10 the difference between FTN and post-training factorization is even more striking in the case of Efficient-Net. For instance, R=1, average accuracy drops to nearly 5%. As R increases, the number of parameters increases sharply but even with 2x more parameters, the average accuracy of post-training low-rank factorization remains lower than FTN with R=1. The Table S9, S11 depicts the approximation error (averaged over all the layers) for different values R and every task. We report average error and standard deviation over the layers for a subset of R. We observe that as R increases approximation error reduces, which leads to improvement in accuracy. Nevertheless, approximation error remains significant even with large R, which is the reason why accuracy remains significantly less than fine-tuned weights. In all cases, accuracy for tasks depends on approximation error (Flowers and CUBS do better than Wikiart and Sketch because the approximation error is small).
@@
Table S8: Low-rank factorization of weight increments on ImageNet-to-Sketch dataset with resnet-50 backbone.
@@
Fine-Tuning(arbitrary $\Delta W_{t}$ = $_{t} -WW$) 6x(141M) 95.69 78.42 81.02 91.44 83.32 85.98
@@
Table S9: Average (std deviation) approximation error for low-rank factorization of weight increments $\Delta W$ on ImageNet-to-Sketch dataset with resnet-50 backbone.
@@
Table S10: Low-rank factorization of weight increments on ImageNet-to-Sketch dataset with efficient-net backbone.
@@
Fine-Tuning (arbitrary $\Delta W_{t}$ = $_{t} -WW$) 6x(105.24M) 96.08 78.72 80.9 92.81 83.67 86.43
@@
Table S11: Average (std deviation) approximation error for low-rank factorization of weight increments $\Delta W$ on ImageNet-to-Sketch dataset with efficient-net backbone.
@@
S6 Multi-domain image generation
@@
Dataset. We used the multi-domain Transient Attributes [20] dataset. The dataset contains a total of 8571 images with 40 annotated attribute labels. Each label is associated with a score in the [1,1][-1,1] range. We utilize the associated confidence score for each season to build our collection of images for each season. Additionally, we only selected images that were captured during daytime. Our training set consisted of $1875$ summer, $1405$ spring, $1353$ autumn, and $2566$ winter images. We normalized each image to the range of [1,1][-1,1] and resized them to a resolution of $128\times 128$.
@@
Figure S7: Generated images for different seasons using FTN.
Results. Table S12 shows the performance, in terms of Peak Signal-to-Noise Ratio (PSNR), for three methods. Our proposed FTN network achieved a comparable performance to the single-domain networks. Each single domain network has 71.471.4M trainable parameters, while the FTN network adds an additional 3.93.9M parameters per domain over the base network. Figure S7 and S8 show examples of images generated by our proposed FTN network. In addition, table S13 shows the average performance of our FTN network under different rank settings. We observe a performance increase by increasing the rank $R$ of our low-rank factors.
Figure S8: Additional generated images for different seasons using FTN.
Table S12: Image generation PSNR for different methods and seasons
@@
Table S13: FTN PSNR for image generation under $R=\{20,50\}$